Find the electric flux through a cylinder with a distance of 22 cm and a height of 35 cm with a charge of 5.42 μC. ( use the general formula of flux and the area of cylinder is 2𝜋rh + 2𝜋r²)



1
Expert's answer
2022-03-13T18:49:05-0400

Answer

Using gauss law flux can be written as

"\phi=\frac{q}{\epsilon_0}\\=\frac{5.42*10^{-6}}{8.85*10^{-12}}\\=0.61*10^{6}Wb"